Income Statement

The accountant for Dolfin Corporation prepared the following income statement. The auditor has asked you to use this statement to provide additional information to Dolfin Corporation on the Final Questions.


Note - If you can, please show/explain how you got the answers. Thank you.
Answer #1

1).

operating expenses.

selling expenses. 41709

administrative expenses. 35292

operating expenses = 41709+35292 = 77001


2).

gross profit.  

gross profit = sales - cost of goods sold

= 320840-218171

= 102669


3). operating income.

operating income = gross profit - operating expenses

= 102669 - 77001

=25668  


answered by: Tata